# Antiferromagnetism and Tightly Bound Cooper Pairs Induced by Kinetic Frustration

### v1

@@ I Introduction
More broadly, a long-standing puzzle persists across families of lightly doped antiferromagnetic insulators. While materials such as LaTiO3, Sr2IrO4, V2O3, NiS2, and Sr2VO4 remain non-superconducting even at ultralow temperatures, others—including electron-doped cuprates [12, 1, 5], heavy-fermion systems like CeRh(Co)In5 [13], and certain organic conductors [14]—exhibit a robust coexistence of antiferromagnetism and superconductivity. This contrast eludes conventional explanations based solely on strong superexchange interactions or low-dimensionality, suggesting that magnetic fluctuations alone are insufficient to account for doping-induced superconductivity.
Identifying a clear pairing mechanism from numerical studies of the doped Hubbard model is notoriously difficult due to the intricate competition between AFM ordering and kinetic energy of doped holes, which leads to multiple intertwined orders. In the strong-coupling regime, superconductivity often competes with robust antiferromagnetic correlations inherited from the Mott insulating state, resulting in a proliferation of nearly degenerate ground states, including stripes, charge density waves, and pair-density wave states [9, 15, 16, 17, 18, 19, 20]. As a result, numerical signatures of superconductivity—such as long-range pairing correlations or anomalous response functions—are often entangled with these other ordering tendencies, complicating efforts to disentangle the intrinsic pairing mechanism. Moreover, finite-size effects and the lack of direct access to dynamical pairing interactions further hinder a definitive interpretation of the results, leaving open the fundamental question of what drives pairing in this paradigmatic model.
Kinetic frustration (KF) has recently emerged as a central mechanism for generating strong pairing from purely repulsive interactions in doped fully polarized Mott insulators [21, 22, 23]. In the strongly correlated regime, frustration of hole motion gives rise to an effective antiferromagnetic interaction between neighboring spins, producing a “counter-Nagaoka effect” in the infinite-$U$ limit [24]. This emergent interaction originates from the formation of a local singlet on triangular plaquettes occupied by a hole, which suppresses KF by inducing an effective $\pi$-flux through the triangle (see inset of Fig. 1). In polarized spin backgrounds, this mechanism leads to the formation of tightly bound hole-magnon composites, or spin polarons [22, 25]. Remarkably, the same kinetic interference responsible for polaron formation can also mediate a strong pairing interaction between polarons, giving rise to a novel superconducting phase termed magnonic superconductivity in regimes of sufficiently strong KF [22, 23].
@@
Here, we leverage the guiding principle of KF to reveal a robust microscopic mechanism that significantly reduces the complexity inherent in the doped Hubbard model. Our central insight is that KF simultaneously stabilizes both antiferromagnetic correlations and superconductivity in the strongly coupling regime, thereby eliminating the primary source of competition that has hindered a clear identification of the pairing mechanism. By combining exact analytical solutions on simplified geometries with large-scale numerical simulations (up to lattices of $16\times 16$ sites) of an extended square-lattice $t$$J$ model with correlated hoppings, we demonstrate that frustration of hole motion—arising from competing nearest-neighbor (NN) and next-nearest-neighbor (NNN) hopping amplitudes, $t_{1}$ and $t_{2}$—gives rise to a superconducting phase coexisting with long-range AFM order. While the resulting pairing mechanism shares similarities with magnonic superconductivity in a ferromagnetic background, a crucial distinction lies in the fact that, in our case, the AFM background is dynamically stabilized by the hole kinetic energy. As shown in Fig. 1, the associated polaronic effect enhances the singlet character of spin pairs on bonds that form triangles with the doped hole to suppress the KF. When a second hole occupies the opposite sublattice, it experiences an effective attraction of order $t_{1}$ when located adjacent to the first hole, driven by the cooperative enhancement of singlet correlations on the two bonds parallel to the two holes. The resulting effective correlated hopping naturally generates strong pairing without requiring any bare attractive interaction.
@@ III Spin fluctuations
Beyond correlation functions, our constrained ED framework provides direct access to the two-hole binding energy - a critical metric for pairing strength. We define the binding gap as the energy difference between the ground state and first excited state within a fixed momentum sector: $\Delta_{g}(k)=E_{1}(k)-E_{0}(k)$. In the thermodynamic limit, the ground state energy $E_{0}(k)$ corresponds to a bound pair while the first excited state energy $E_{1}(k)$ corresponds to the lowest unpaired state in the scattering continuum. Thus, $\Delta_{g}(k)$ serves as a direct measure of the superconducting pairing gap.
To minimize finite-size effects, we implement twisted boundary conditions by choosing momentum $k=(0,2\pi/L)$. This choice imposes distinct boundary conditions on the relative motion of the two-hole system: periodic along one spatial direction and antiperiodic along the orthogonal direction. Critically, this configuration suppresses spurious finite-size effects—specifically, the artificial binding gap inherent to non-interacting systems under conventional periodic PBC—thereby isolating interaction-driven pairing contributions (see SM for analytical justification). In the thermodynamic limit ($L\to\infty$), the $\mathbf{k}=(0,2\pi/L)$ sector converges to $\mathbf{k}=0$, where the Cooper pair ground state naturally resides.
@@ IV AFM polaron enhanced hole pairing
To quantify how singlet formation enhances kinetic energy, we decompose the total kinetic energy across different bonds. The kinetic energy $H_{t}=-\sum_{jk,\sigma}t_{jk}c_{j\sigma}^{\dagger}c_{k\sigma}$ of the two-hole system can be expressed as the expectation value of correlated hopping processes where one hole at site $j$ moves to a neighboring site, while the other hole remains fixed at a given site $l$:
@@
Figure 6: (a) Modified hopping amplitude $\tilde{t}_{jk}(l)$ on NN bonds calculated using U(1) DMRG on an 8-leg cylinder with up to 8000 states per block and parameters $\lambda=\Delta=1,\,U=10t_{1}$. (b) Modified hopping amplitude on NNN bonds on opposite sublattice of fixed holes using identical computational parameters.
Fig. 6(a, b) shows the effective hopping amplitudes $\tilde{t}_{jk}(l)$ computed with DMRG for two holes on an 8-leg cylinder of length $L_{x}=16$. As anticipated from Fig. 5 and the suppression of kinetic frustration via enhanced singlet formation on the triangle formed by the hole and two neighboring spins, $\tilde{t}_{jk}(l)$ reaches its maximum when the hole at site $l$ lies adjacent to the bond connecting sites $j$ and $k$. In other words, as illustrated in Fig. 5(c), strong singlet correlations on the bonds parallel to the two holes suppress kinetic frustration and enhance $\tilde{t}_{jk}(l)$ when $j$, $k$, and $l$ form a minimal triangle.
@@
To verify that the polaronic mechanism—where spin fluctuations suppress kinetic frustration—provides the dominant contribution to the Cooper pair binding energy, we solved an effective two-particle problem using the correlated hopping amplitudes $\tilde{t}_{jk}(l)$ extracted from our DMRG calculations on the 8-leg cylindrical cluster shown in Fig. 6. We assumed that these amplitudes remain unchanged beyond the largest distances accessible in the cluster (see Supplementary Material). In the thermodynamic limit, this yields a coherence length of $\xi=9.6$, which is reduced to $\xi=4.6$ when including the nearest-neighbor attractive density-density interaction arising from the Ising spin-spin interaction discussed in Section II. This result is in good agreement with the coherence length $\xi\simeq 6.3$ obtained from our truncated ED calculations on the largest accessible $16\times 16$ system, using the same Hamiltonian parameters $(\lambda=\Delta=1,\,U=10t_{1})$. In other words, the attractive interaction generated by the polaronic mechanism through correlated effective hopping is sufficient to explain a coherence length of only a few lattice spacings.
@@ V Avoidance of phase separation
Having established both strong pairing and phase separation suppression, we estimate superconducting transition temperatures using our calculated binding gaps.For sufficiently low hole concentration $\rho_{h}$, such that $\xi\rho_{h}^{1/2}\ll 1$, the system enters the Bose-Einstein condensation (BEC) regime, where superconductivity arises from the superfluid transition of preformed pairs. In this limit, the critical temperature depends only on the superfluid density $\rho\simeq\rho_{h}/2$ and the effective mass of the pairs $m_{b}$, and is given by $k_{B}T_{c}\approx C\frac{\hbar^{2}\rho}{m_{b}}$, where $C\approx\frac{2\pi}{\log(380)}$ is the contact repulsive pair-pair interaction in the continuum (long wavelength) limit. Hole concentration level and weak interactions modify the constant $C$ only slightly though a double logarithmic dependence [43]. For the full $t$-$J$ model with $(\lambda,\Delta)=(1,1)$, our spectrum for a $16\times 16$ cluster with periodic PBC yields $k_{B}T_{c}\approx 1.16\,\rho_{h}\,t_{1}$ in the BEC regime.
@@ VI Conclusions
In essence, the “bipolaron” formed around a hole pair enables effective correlated hopping through multi-$\pi$-flux generation, which is strong enough to bind the two holes. This mechanism is, in some sense, analogous to the spin-bag scenario proposed by Schrieffer for the unfrustrated case [47], where kinetic energy promotes ferromagnetic correlations in accordance with Nagaoka’s theorem [48]. A key difference, however, lies in the energy scales: the spin-bag mechanism requires very large values of $U/t_{1}$ for polarons to form with appreciable size, since ferromagnetic polarization is strongly penalized by the antiferromagnetic exchange $J_{1}$. In contrast, the antiferromagnetic polaron emerging in our model has a relatively large spatial extent even for realistic values of $U/t_{1}\simeq 10$, simply because both the kinetic energy and $J_{1}$ cooperatively favor singlet formation. The polaron arises naturally because $J_{1}$ favors a uniform distribution of singlet character over all bonds, while the kinetic energy drives an enhancement of singlet correlations in the vicinity of each hole. While kinetic frustration accounts for the majority of pairing enhancement, the AFM exchange coupling $J_{1}$ provides the remaining contribution.
@@ A.1 Sublattice ferromagnetism through Nagaoka mechanism at $t_{1}=0$
$|\Psi_{G}\rangle=\sum_{r_{1},r_{2}}\psi(r_{1},r_{2})|\phi(r_{1},r_{2})\rangle$ (1)
where $\psi(r_{1},r_{2})$ are superposition coefficients for configurations with holes at positions $r_{1}$ and $r_{2}$, and $|\phi(r_{1},r_{2})\rangle$ denotes a many-body state with holons at sites $r_{1},r_{2}$ and arbitrary spin configurations on the remaining sites. These states satisfy the orthonormality condition $\langle\phi(r_{1},r_{2})|\phi(r_{1}^{\prime},r_{2}^{\prime})\rangle=\delta_{r_{1},r_{1}^{\prime}}\delta_{r_{2},r_{2}^{\prime}}$, and the wavefunction is normalized such that $\langle\Psi_{G}|\Psi_{G}\rangle=1$.
@@
$E_{0}=\langle\Psi_{G}|\mathcal{H}|\Psi_{G}\rangle$$\geq$ $-2t_{2}\sum_{r_{1},\langle\langle r_{2},r_{3}\rangle\rangle}|\psi(r_{1},r_{2})^{*}\psi(r_{1},r_{3})|$ (2)
$\geq$ $-4t_{2}\sum_{r_{1},r_{2}}|\psi(r_{1},r_{2})|^{2}=-8t_{2}$
@@
$\psi(r_{1},r_{2})=\frac{2}{N}(-1)^{(x_{1}+x_{2}-1)}$ (3)
for $r_{1},r_{2}$ located on different sublattices, and zero otherwise. Here, $x_{1}$ and $x_{2}$ denote the $x$-coordinates of $r_{1}$ and $r_{2}$, respectively. The optimal spin background features uniform spin alignment within each sublattice, while the relative orientation between the two sublattices remains arbitrary:
$|\phi(r_{1},r_{2})\rangle=c_{\hat{n}_{1},r_{1}}c_{\hat{n}_{2},r_{2}}|\hat{n}_{1}\hat{n}_{2}\rangle$ (4)
@@
$|\hat{n}_{1}\hat{n}_{2}\rangle=\prod_{i\in A}c^{\dagger}_{\hat{n}_{1},i}\prod_{j\in B}c^{\dagger}_{\hat{n}_{2},j}|0\rangle$ (5)
Here, $c^{\dagger}_{\hat{n},i}$ creates a fermion at site $i$ with spin pointing along unit vector $\hat{n}$. The state $|\hat{n}_{1}\hat{n}_{2}\rangle$ corresponds to a spin configuration in which all spins on sublattice A align along $\hat{n}_{1}$, while those on sublattice B align along $\hat{n}_{2}$.

### v2

@@ I Introduction
More broadly, a long-standing puzzle persists across families of lightly doped antiferromagnetic insulators. While materials such as LaTiO3, Sr2IrO4, V2O3, NiS2, and Sr2VO4 remain non-superconducting even at ultralow temperatures, others—including electron-doped cuprates [12, 1, 13, 5], heavy-fermion systems like CeRh(Co)In5 [14], and certain organic conductors [15]—exhibit a robust coexistence of antiferromagnetism and superconductivity. This contrast eludes conventional explanations based solely on strong superexchange interactions or low-dimensionality, suggesting that magnetic fluctuations alone are insufficient to account for doping-induced superconductivity.
Identifying a clear pairing mechanism from numerical studies of the doped Hubbard model is notoriously difficult due to the intricate competition between AFM ordering and kinetic energy of doped holes, which leads to multiple intertwined orders. In the strong-coupling regime, superconductivity often competes with robust antiferromagnetic correlations inherited from the Mott insulating state, resulting in a proliferation of nearly degenerate ground states, including stripes, charge density waves, and pair-density wave states [16, 9, 17, 18, 19, 20, 21, 22]. As a result, numerical signatures of superconductivity—such as long-range pairing correlations or anomalous response functions—are often entangled with these other ordering tendencies, complicating efforts to disentangle the intrinsic pairing mechanism. Moreover, finite-size effects and the lack of direct access to dynamical pairing interactions further hinder a definitive interpretation of the results, leaving open the fundamental question of what drives pairing in this paradigmatic model.
Kinetic frustration (KF) has recently emerged as a central mechanism for generating strong pairing from purely repulsive interactions in doped fully polarized Mott insulators [23, 24, 25]. In the strongly correlated regime, frustration of hole motion gives rise to an effective antiferromagnetic interaction between neighboring spins, producing a “counter-Nagaoka effect” in the infinite-$U$ limit [26, 27, 28, 29]. This emergent interaction originates from the formation of a local singlet on triangular plaquettes occupied by a hole, which suppresses KF by inducing an effective $\pi$-flux through the triangle (see inset of Fig. 1). In polarized spin backgrounds, this mechanism leads to the formation of tightly bound hole-magnon composites, or spin polarons [24, 30]. Remarkably, the same kinetic interference responsible for polaron formation can also mediate a strong pairing interaction between polarons, giving rise to a novel superconducting phase termed magnonic superconductivity in regimes of sufficiently strong KF [24, 25].
@@
Here, we leverage the guiding principle of KF to reveal a robust microscopic mechanism that significantly reduces the complexity inherent in the doped Hubbard model. Our central insight is that KF simultaneously stabilizes both antiferromagnetic correlations and superconductivity in the strongly coupling regime, thereby eliminating the primary source of competition that has hindered a clear identification of the pairing mechanism. By combining exact analytical solutions on simplified geometries with large-scale numerical simulations (up to lattices of $16\times 16$ sites) of an extended square-lattice $t$$J$ model with correlated hoppings, we demonstrate that frustration of hole motion—arising from competing nearest-neighbor (NN) and next-nearest-neighbor (NNN) hopping amplitudes, $t_{1}$ and $t_{2}$—gives rise to a superconducting phase coexisting with long-range AFM order. While the resulting pairing mechanism shares similarities with magnonic superconductivity in a ferromagnetic background, a crucial distinction lies in the fact that, in our case, the AFM background is dynamically stabilized by the hole kinetic energy. As shown in Fig. 1, the associated polaronic effect enhances the singlet character of spin pairs on bonds that form triangles with the doped hole to suppress the KF. When a second hole occupies the opposite sublattice, it experiences an effective attraction of order $t_{1}$ when located adjacent to the first hole, driven by the cooperative enhancement of singlet correlations on the two bonds parallel to the two holes. The resulting two-body hopping naturally generates strong pairing without requiring any bare attractive interaction.
@@ III Spin fluctuations
Beyond correlation functions, our constrained ED framework provides direct access to the two-hole binding energy - a critical metric for pairing strength. We define the binding gap as the energy difference between the ground state and first excited state within a fixed momentum sector: $\Delta_{g}(\bm{k})=E_{1}(\bm{k})-E_{0}(\bm{k})$. In the thermodynamic limit, the ground state energy $E_{0}(\bm{k})$ corresponds to a bound pair while the first excited state energy $E_{1}(\bm{k})$ corresponds to the lowest unpaired state in the scattering continuum. Thus, $\Delta_{g}(\bm{k})$ serves as a direct measure of the superconducting pairing gap.
To minimize finite-size effects, we implement twisted boundary conditions by choosing momentum $k=(0,2\pi/L)$. This choice imposes distinct boundary conditions on the relative motion of the two-hole system: periodic along one spatial direction and antiperiodic along the orthogonal direction. Critically, this configuration suppresses spurious finite-size effects—specifically, the artificial binding gap inherent to non-interacting systems under conventional periodic PBC—thereby isolating interaction-driven pairing contributions (see SM for analytical justification). In the thermodynamic limit ($L\to\infty$), the $\bm{k}=(0,2\pi/L)$ sector converges to $\bm{k}=0$, where the Cooper pair ground state naturally resides.
@@ IV AFM polaron enhanced hole pairing
To quantify how singlet formation enhances kinetic energy, we decompose the total kinetic energy across different bonds. The kinetic energy $H_{t}=-\sum_{jk,\sigma}t_{jk}c_{j\sigma}^{\dagger}c_{k\sigma}$ of the two-hole system can be expressed as the expectation value of two-body hopping processes where one hole at site $j$ moves to a neighboring site, while the other hole remains fixed at a given site $l$:
@@
Figure 6: (a) two body hopping amplitude $\tilde{t}_{jk}(l)$ on NN bonds calculated using U(1) DMRG on an 8-leg cylinder with up to 8000 states per block and parameters $\lambda=\Delta=1,\,U=10t_{1}$. (b) Modified hopping amplitude on NNN bonds on opposite sublattice of fixed holes using identical computational parameters.
Fig. 6(a, b) shows the effective two-body hopping amplitudes $\tilde{t}_{jk}(l)$ computed with DMRG for two holes on an 8-leg cylinder of length $L_{x}=16$. As anticipated from Fig. 5 and the suppression of kinetic frustration via enhanced singlet formation on the triangle formed by the hole and two neighboring spins, $\tilde{t}_{jk}(l)$ reaches its maximum when the hole at site $l$ lies adjacent to the bond connecting sites $j$ and $k$. In other words, as illustrated in Fig. 5(c), strong singlet correlations on the bonds parallel to the two holes suppress kinetic frustration and enhance $\tilde{t}_{jk}(l)$ when $j$, $k$, and $l$ form a minimal triangle.
@@
To verify that the polaronic mechanism—where spin fluctuations suppress kinetic frustration—provides the dominant contribution to the Cooper pair binding energy, we solved an effective two-particle problem using the two-body hopping amplitudes $\tilde{t}_{jk}(l)$ extracted from our DMRG calculations on the 8-leg cylindrical cluster shown in Fig. 6. We assumed that these amplitudes remain unchanged beyond the largest distances accessible in the cluster (see Supplementary Material). In the thermodynamic limit, this yields a coherence length of $\xi=9.6$, which is reduced to $\xi=4.6$ when including the nearest-neighbor attractive density-density interaction arising from the Ising spin-spin interaction discussed in Section II. This result is in good agreement with the coherence length $\xi\simeq 6.3$ obtained from our truncated ED calculations on the largest accessible $16\times 16$ system, using the same Hamiltonian parameters $(\lambda=\Delta=1,\,U=10t_{1})$. In other words, the attractive interaction generated by the polaronic mechanism through two-body hopping is sufficient to explain a coherence length of only a few lattice spacings.
@@ V Avoidance of phase separation
Having established both strong pairing and phase separation suppression, we estimate superconducting transition temperatures using our calculated binding gaps.For sufficiently low hole concentration $\rho_{h}$, such that $\xi\rho_{h}^{1/2}\ll 1$, the system enters the Bose-Einstein condensation (BEC) regime, where superconductivity arises from the superfluid transition of preformed pairs. In this limit, the critical temperature depends only on the superfluid density $\rho\simeq\rho_{h}/2$ and the effective mass of the pairs $m_{b}$, and is given by $k_{B}T_{c}\approx C\frac{\hbar^{2}\rho}{m_{b}}$, where $C\approx\frac{2\pi}{\log(380)}$ is the contact repulsive pair-pair interaction in the continuum (long wavelength) limit. Hole concentration level and weak interactions modify the constant $C$ only slightly though a double logarithmic dependence [48]. For the full $t$-$J$ model with $(\lambda,\Delta)=(1,1)$, our spectrum for a $16\times 16$ cluster with PBC yields $k_{B}T_{c}\approx 1.16\,\rho_{h}\,t_{1}$ in the BEC regime.
@@ VI Conclusions
In essence, the “bipolaron” formed around a hole pair enables effective two-body hopping through multi-$\pi$-flux generation, which is strong enough to bind the two holes. This mechanism is, in some sense, analogous to the spin-bag scenario proposed by Schrieffer for the unfrustrated case [42, 52], where kinetic energy promotes ferromagnetic correlations in accordance with Nagaoka’s theorem [36]. A key difference, however, lies in the energy scales: the spin-bag mechanism requires very large values of $U/t_{1}$ for polarons to form with appreciable size, since ferromagnetic polarization is strongly penalized by the antiferromagnetic exchange $J_{1}$. In contrast, the antiferromagnetic polaron emerging in our model has a relatively large spatial extent even for realistic values of $U/t_{1}\simeq 10$, simply because both the kinetic energy and $J_{1}$ cooperatively favor singlet formation. The polaron arises naturally because $J_{1}$ favors a uniform distribution of singlet character over all bonds, while the kinetic energy drives an enhancement of singlet correlations in the vicinity of each hole. While kinetic frustration accounts for the majority of pairing enhancement, the AFM exchange coupling $J_{1}$ provides the remaining contribution.
@@ A.1 Sublattice ferromagnetism through Nagaoka mechanism at $t_{1}=0$
$|\Psi_{G}\rangle=\sum_{\bm{r}_{1},\bm{r}_{2}}\psi(\bm{r}_{1},\bm{r}_{2})|\phi(\bm{r}_{1},\bm{r}_{2})\rangle$ (1)
where $\psi(\bm{r}_{1},\bm{r}_{2})$ are superposition coefficients for configurations with holes at positions $\bm{r}_{1}$ and $\bm{r}_{2}$, and $|\phi(\bm{r}_{1},\bm{r}_{2})\rangle$ denotes a many-body state with holons at sites $\bm{r}_{1},\bm{r}_{2}$ and arbitrary spin configurations on the remaining sites. These states satisfy the orthonormality condition $\langle\phi(\bm{r}_{1},\bm{r}_{2})|\phi(\bm{r}_{1}^{\prime},\bm{r}_{2}^{\prime})\rangle=\delta_{\bm{r}_{1},\bm{r}_{1}^{\prime}}\delta_{\bm{r}_{2},\bm{r}_{2}^{\prime}}$, and the wavefunction is normalized such that $\langle\Psi_{G}|\Psi_{G}\rangle=1$.
@@
$E_{0}=\langle\Psi_{G}|\mathcal{H}|\Psi_{G}\rangle$$\geq$ $-2t_{2}\sum_{\bm{r}_{1},\langle\langle\bm{r}_{2},\bm{r}_{3}\rangle\rangle}|\psi(\bm{r}_{1},\bm{r}_{2})^{*}\psi(\bm{r}_{1},\bm{r}_{3})|$ (2)
$\geq$ $-4t_{2}\sum_{\bm{r}_{1},\bm{r}_{2}}|\psi(\bm{r}_{1},\bm{r}_{2})|^{2}=-8t_{2}$
@@
$\psi(\bm{r}_{1},\bm{r}_{2})=\frac{2}{N}(-1)^{(x_{1}+x_{2}-1)}$ (3)
for $\bm{r}_{1},\bm{r}_{2}$ located on different sublattices, and zero otherwise. Here, $x_{1}$ and $x_{2}$ denote the $x$-coordinates of $\bm{r}_{1}$ and $\bm{r}_{2}$, respectively. The optimal spin background features uniform spin alignment within each sublattice, while the relative orientation between the two sublattices remains arbitrary:
$|\phi(\bm{r}_{1},\bm{r}_{2})\rangle=c_{\hat{\bm{n}}_{1},\bm{r}_{1}}c_{\hat{\bm{n}}_{2},\bm{r}_{2}}|\hat{\bm{n}}_{1}\hat{\bm{n}}_{2}\rangle$ (4)
@@
$|\hat{\bm{n}}_{1}\hat{\bm{n}}_{2}\rangle=\prod_{i\in A}c^{\dagger}_{\hat{\bm{n}}_{1},i}\prod_{j\in B}c^{\dagger}_{\hat{\bm{n}}_{2},j}|0\rangle$ (5)
Here, $c^{\dagger}_{\hat{\bm{n}},i}$ creates a fermion at site $i$ with spin pointing along unit vector $\hat{\bm{n}}$. The state $|\hat{\bm{n}}_{1}\hat{\bm{n}}_{2}\rangle$ corresponds to a spin configuration in which all spins on sublattice A align along $\hat{\bm{n}}_{1}$, while those on sublattice B align along $\hat{\bm{n}}_{2}$.